Question A

When a company participates in a stock buyback program, it means that the company is buying shares of its own stock and taking them off the market. With this simple definition in mind, how would a company's stock buyback program affect its Earnings per Share?
Answer #1

When a company buys back its own shares, the number of shares reduce. As a result, the EPS will increase ( since EPS = Net Income/Number of shares).

Assuming a company has $ 10,000,000 as its net income and the outstanding shares are 1,000,000 numbers. The EPS = 10,000,000/1000,000 = 10$.

Now if the company buys back 250,000 shares from open market, the number of shares outstanding will be 750,000 numbers. So, EPS = 10,000,000/750,000 = 13.33 $. Hence EPS rises by 33 1/3 %.